The amount of tax is directly proportional to the cost of an item. If the tax on a $12,300 car is $738, find the tax on a $9,450 car.
Answer by sarah_adam(201)

Given Tax is directly proportional to cost
T = tax and C0st = C
T1*C1 = T2*C2
T1 = $738, C1 = $12,300 , T2 , C2 = $9450
738*12,300 = T2 * 9450
T2 = (738*12,300)/9450 = $960.57